Solve the equation; x2 - 8x + 7 = 0 using the Quadratic Formula


1
Expert's answer
2020-09-10T18:12:20-0400

Solutions of the quadratic equation

. "ax^{2}+bx+c=0,\; a\neq0\;\;are-
\\
\\
x =\dfrac{-b \plusmn \sqrt{b^2-4ac}}{2a}
\\
\\
Here, we\:need\:to\:solve-
\\
x^{2}-8x+7=0
\\
i.e., a=1, b=-8,c=7 \\
so, x=\dfrac{-(-8)\plusmn\sqrt{(-8)^{2}-4(1)(7)}}{2(1)}
\\ 
\implies x = \dfrac{8 \plusmn \sqrt{64-28}}{2}"

"\implies x = \dfrac{8 \plusmn 6}{2} =7,1"